OpenStudy (anonymous):

Find a1 for the given geometric series Sn= 89205 R=4 N=11 ANSWERS: A: 0.11 B: 0.06 C. 8089.55 D: 60.93

OpenStudy (goformit100):

Where is the equation ?

OpenStudy (anonymous):

Do you know the equation of the the sum of a geometric series?

OpenStudy (anonymous):

no the question only says what i wrote above ^ i'm really confused

OpenStudy (anonymous):

Ok, well the equation for a geometri series (a series is a sum of numbers ina sequece and a sequence is just a list of numbers) is \[\Huge S _{n}=a _{1}(\frac{1-r^n}{1-r})\]

OpenStudy (anonymous):

Sn=the sum of the numbers a1=the first term of the series, and what we're trying to find r=the common ratio, which is the number that you multiply the previus number to n=the number of terms Well we have Sn, r, and n, so all we have to have to is substitute and find a1 which in this is our x

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

here r=4 >1 so Sn = a1. (r^n-1) /(r-1) or a1=((r-1)*Sn)/( (r^n-1)

OpenStudy (anonymous):

True^ @matricked

OpenStudy (anonymous):

here r=4 ,n=11

OpenStudy (jhannybean):

because r =4> 1 (diverges) that classifies S_n = a1? How so...just wondering.

OpenStudy (anonymous):

check whether ans u have posted are of the same question

OpenStudy (jhannybean):

I was just referring to your post "matricked here r=4 >1 so Sn = a1. (r^n-1) /(r-1) or a1=((r-1)*Sn)/( (r^n-1)"

OpenStudy (anonymous):

the question and answers i posted are from the same question. those are the selected answers

OpenStudy (anonymous):

Well you've been gien the equation \[\Huge a_1=\frac{ (r-1)S_n }{ r^n-1 }\]

OpenStudy (anonymous):

And you have r,n,and Sn

OpenStudy (anonymous):

\[a _{1}= (3)(89205)\div 4^{10}\]

OpenStudy (anonymous):

a bit change a1= ((3)(89205))÷(4^11 -1) = 0.0638 appx =0.06

OpenStudy (anonymous):

thank you i get it now

OpenStudy (anonymous):

yw